OpenStudy (anonymous):

What is the angle of refraction when a ray of light passes from ethanol to air if the incident angle is 34°? The refractive index of ethanol is 1.36 and of air is 1.00. 21° 34° 49° 66°

OpenStudy (danjs):

Snell's Law: The ratio of the sines of the angles of incidence and refraction is equal to the reciprocal of the ratio of indices of refraction. \[\frac{ Sin(\theta _{E} )}{ Sin(\theta _{A})}=\frac{ n _{A} }{ n _{E} }\] 3 of the 4 values are given.